OpenStudy (anonymous):

a+6a-9=30=1a+6a-9=30=7a-9=30=7a÷39=A=5.5714285 (Right?)

OpenStudy (lgbasallote):

uhh might be a good idea to write it in a more organized way? are they really in one line only?

OpenStudy (anonymous):

a+6a-9=30 1a+6a-9=30 7a-9=30 7a÷39=5.5714285 A=5.5714285

OpenStudy (lgbasallote):

what you did on the 4th line was weird...but you're right

OpenStudy (lgbasallote):

i think the 4th line should read 7a÷7=5.5714285

OpenStudy (lgbasallote):

i assume it was a typo

OpenStudy (anonymous):

how can you make the answer smaller???

OpenStudy (lgbasallote):

round it off?

OpenStudy (lgbasallote):

usually i go for 5.6 or 5.57